### v1

@@ 6 Conclusion
In this work, we have introduced a comprehensive framework for semantic change detection in conflict zones through three key contributions. First, we presented a novel dataset specifically designed for assessing conflict-induced land changes, featuring six critical damage and construction categories. Second, we proposed the CSD paradigm, which directly focuses on changed areas to enable efficient damage assessment while significantly reducing annotation costs. Third, we developed the MC-DiSNet that effectively leverages pre-trained vision foundation models to enhance feature representation for small targets. Our experimental results demonstrate that the proposed method achieves strong performance on both Gaza-change and SECOND datasets.
This work takes a significant step toward automated conflict damage assessment using remote sensing imagery. We believe our contributions in dataset creation, task formulation, and methodology development will facilitate future research in humanitarian remote sensing and emergency response.

### v2

@@ 6 Conclusion
In this work, we have introduced a comprehensive framework for semantic change detection in conflict zones through three key contributions. First, we presented a novel dataset specifically designed for assessing conflict-induced land changes, featuring six critical damage and construction categories. Second, we introduced the MCD paradigm, which directly focuses on changed areas to enable efficient damage assessment while significantly reducing annotation costs. Third, we developed the MC-DiSNet that effectively leverages pre-trained vision foundation models to enhance feature representation for small targets. Our experimental results demonstrate that the proposed method achieves strong performance on the Gaza-change, SECOND, and Landsat-SCD datasets. This work takes a significant step toward automated conflict damage assessment using remote sensing imagery. We believe our contributions in dataset creation, task formulation, and methodology development will facilitate future research in humanitarian remote sensing and emergency response.